OpenStudy (anonymous):

simplify : ^3√56x^14

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt[3]{56x^{14}}\] \[\LARGE \sqrt[3]{8*7x^{12}*x^2}\] \[\LARGE \sqrt[3]{8}*\sqrt[3]{7}*\sqrt[3]{x^{12}}*\sqrt[3]{x^2}\] \[\LARGE 2*\sqrt[3]{7}*x^4*\sqrt[3]{x^2}\] \[\LARGE 2x^4\sqrt[3]{7x^2}\] So \[\LARGE \sqrt[3]{56x^{14}}=2x^4\sqrt[3]{7x^2}\]